Please Help I can not get this problem.
A triangular region of a parking lot was measured. The measure of the first angle was twice the measure of the second angle. The measure of the third angle is five more than the measure of the first angle. What are the measurements of all three angles? And how do I check my answer.

Found 2 solutions by mathmaven53, DWL:
Answer by mathmaven53(29)

Let second angle be x
Then first angle is 2x
Third angle is 2x + 5
The sum of the angles of a triangle is 180 degrees
x + 2x + 2x + 5 = 180
Now solve for x
Then first angle is 2x
second angle is x
third angle is 2x + 5

Answer by DWL(56)

You need to know that the sum of all the angles in a triangle = 180.
so in your problem:
you can write your problem 2a + a + 5a = 180
combining the "a" we get 8a=180
divide out the 8 on both sideS: a = 180 / 8
a = 22.50
answer:
the second angle = 22.50
The first angle = 2 x 22.50 = 45.0
The third angle = 5 x 22.50 = 112.50
to check, add all the angles AND THEY MUST EQUAL 180 = 22.50 + 45.0 + 112.50 = 180.